# Perturbation-based Analysis of Compositional Data

### v1

@@ 4.3 Variable influence measures when predicting BMI from gut microbiome
While the average perturbation effect estimates which are specifically tailored to the compositional setting are consistent across different regression methods, the permutation and $R^{2}$ measures are – unsurprisingly – not. The population version of the nonparametric $R^{2}$ is always 00, so the estimates are essentially mean zero random noise that depends on the estimator. For the permutation measure the inconsistency comes the fact that the permuted data inherently leave the support of the simplex, making the resulting influence measure highly dependent on how the regression methods extrapolate. Unfortunately, due to a lack of alternatives, it is still common to use these types feature importance or influence measures in microbiome science (e.g., Marcos-Zambrano et al., 2021). Furthermore, the results show that the nonparametric estimator of the CKE is less well-behaved than the other estimators based on the partially linear model. One reason for this could be that dividing by the propensity score estimator results in a more unstable estimate, particularly when the propensity score is poorly estimated.
Next, we empirically demonstrate that simply ignoring the simplex constraint is not an option. To this end we consider a standard approach for estimating effects of coordinates for unconstrained Euclidean data. More specifically, we compare the proposed $\theta_{j}=\mathrm{CFI}_{\mathrm{unit}}^{j}$ (with estimator based on the partially linear model) with a similar estimator but based on the partially linear model $\mathbb{E}[Y\mid Z]=\theta_{j}Z^{j}+h(Z^{-j})$ in which the parameter $\theta_{j}$ is unidentified due to the simplex constraint. For both methods we use a regression estimator that uses cross-validation to select the best estimator out of a linear regression, random forest and support vector regression. As expected the results from the estimator that ignores the simplex structure are meaningless; for example, $\mathrm{var}(Z^{j})\widehat{\theta}_{j}^{2}/\mathrm{var}(Y)$, a proxy for the variance explained by the linear effect, is larger than $10^{17}$ for all species. In contrast, for $\mathrm{CFI}_{\mathrm{unit}}$ these values lie between $10^{-7}$ and $1.6$.
@@ 6 Discussion
We introduced perturbations on the simplex as a formalization of hypothetical changes to a compositional vector and used them to construct interpretable target parameters, which we call average perturbation effects. In our view, perturbations provide a natural way for practitioners to formalize research questions about non-standard changes to a system. We illustrated this, for example, by using perturbations to formalize the otherwise vague concept of an effect of diversity as well as to construct nonparametric feature influence measures for compositional coordinates. Unfortunately, many applied researchers working with compositional data, generally due to a lack of alternatives, still rely on marginal effects by analyzing dependence between a summary statistic and a response (e.g., by regressing a response on a diversity measure). As we demonstrated, such approaches can be misleading since they do not account for the confounding induced by the remaining variation in compositional vector. We hope that this work can draw attention to such pitfalls and improve the analysis of such data by providing a set of easily accessible alternative methods.
@@
We additionally proposed efficient estimators for the average perturbation effects based on semiparametric theory and provided theoretical guarantees that enable us to perform valid inference. Most statistical methods for compositional data analysis to date are based on the parametric log-ratio approach introduced in the seminal paper Aitchison (1982). Our nonparametric approach goes beyond these established methods and allows practitioners to apply theoretically justified procedures to modern compositional data applications. Importantly, our target parameters and estimators circumvent the zero problem central to any log-ratio based analysis and naturally apply to high-dimensional data. This makes our proposed estimators applicable to a range of modern applications that previously had to rely on ad hoc methods or simply ignore the compositional constraints. An additional advantage of a perturbation-based analysis is that it easily facilitates a causal rather than observational interpretation, allowing practitioners to make standard causal inference assumptions to derive rigorous causal conclusions.
@@
In this work we have focused on data constraints induced by the simplex. There are however many other types of constraints that can arise, e.g., directional data, general manifold-valued data or otherwise constrained Euclidean data. The idea of defining target parameters via perturbations easily extends to those settings. Similarly, perturbations can also be used to define meaningful effects of summary statistics on unconstrained data; something which has received surprisingly little attention in the literature but relates to problems in causal inference(e.g., how to define causal variables). Nevertheless, for both extensions, more work is required to develop reparametrizations that can be used to derive efficient estimators akin to ours using semiparametric theory. We hope our work will spark new developments in this area and help provide practitioners with the statistical tools they need.
@@ Acknowledgements
We thank Alexander Mangulad Christgau and Lucas Kook for helpful discussions. ARL and NP were supported by a research grant (0069071) from Novo Nordisk Fonden.
@@
(b) By the same argument as in (a), we have that $\phi$ is invertible with inverse as computed in (S3). This reparametrization is derivative-isolating for perturbations with direction $v$ given by (6) and speed given by (reversing the computation in (a))
$s:z\mapsto-\partial_{\ell}t_{(E(z),v(z))}^{-1}(t(z))=-(\partial_{\delta}t_{(E(z),v(z))}(\|E(z)-z\|_{1}))^{-1}.$
@@
$L:=W^{1}+(1+\mathbbm{1}_{\{W^{1}>1/2\}})\xi,\quad Y:=(1+L)W^{1}+\varepsilon.$
@@
$\min_{\theta\in\mathbb{R},\widetilde{h}:\mathcal{W}\to\mathbb{R}}\underbrace{\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W]-\theta(L-\mathbb{E}[Y\,|\,W])-\widetilde{h}(W)\right)^{2}\right]}_{=c(\theta,\widetilde{h})},$ (S5)
@@
2:For $i\in[n]$, define $D_{i}:=\mathbbm{1}_{\{\omega_{\psi}(Z_{i})\neq 0\}}$.
@@
Section S4.1 provides the details behind the experiments in Table 1 in Section 1. Section S4.2 contains additional simulation results from our investigation into the performance of various semiparametric estimators. Section S4.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 4.2.
@@
$Y\,|\,Z\sim\mathrm{Bern}\left(\frac{3}{4}-\frac{1}{8}\mathbbm{1}_{\{Z^{1}=0\}}+\frac{1}{2}\mathbbm{1}_{\{Z^{2}=0\}}\right).$
@@
S4.3 Description of the construction of groups in Section 4.2
In this section we describe how we construct the two community-level grouped datasets analyzed in Section 4.2. On the individual-level data the observed proportions of race (categorical, ’White’, ’Black’ and ’Other’), education (binary, $1$ if strictly greater than high school and 00 otherwise) and compensation (binary, $1$ if compensation above 50,000 USD per year and 00 otherwise) are
@@
S4.4 Additional plots from gut microbiome analysis in Section 4.3
@@
$\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon)\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(\{|(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon\}\cap\Omega_{P})+\zeta/2\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon)+\zeta/2,$
hence it suffices to prove that $\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ to conclude that $(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$. To that end, we first note that
$\mathbb{E}_{P}\left[\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})\,|\,\widehat{f}_{k}\right]=\mathbb{E}_{P}\left[\mathbbm{1}_{\Omega_{P}}\left(\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}\right)\,|\,\widehat{f}_{k}\right].$
@@
$\mathbb{E}_{P}\left[\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})\,|\,\widehat{f}_{k}=h\right]=0$ (S11)
for $\mathbb{P}_{P}$-almost every $h$. If $h$ is such that $\mathbbm{1}_{\Omega_{P}}=0$ then clearly the conditional expectation is 00. If $h$ is such that $\mathbbm{1}_{\Omega_{P}}=1$, then we intend to apply Lemma S5 (integration by parts) with $g=(h-f_{P})$. Assumptions (a)-(c) of Lemma S5 are clear from the above and (d) follows from Lemma S6 and Assumption (a). We conclude that
@@
$\mathbb{P}_{P}(|\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon\,|\,\widehat{f}_{k})$
$\qquad\leq\frac{1}{\epsilon^{2}N_{k,r}}\mathbb{E}_{P}[\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})^{2}\,|\,\widehat{f}_{k}]$
@@
$\mathbb{P}_{P}(|\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\,|\,\widehat{f}_{k})=o_{\mathcal{P}}(1)$
which implies that $\mathbbm{1}_{\Omega_{P}}\cdot(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ so $(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ by our earlier argument. The above combined with Lemma S1 proves (S10).
@@
and (S12) implies that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(S_{i}-\mathrm{CFI}^{j})=O_{\mathcal{P}}(1)$, we have by Lemma S3, (S10) and (S13) that
@@
$\mathrm{var}_{P}(S_{i})$ $=\mathbb{E}_{P}\left[\left(\partial_{\ell}\widehat{f}(L,W)-\rho_{P}(L,W)(Y-f_{P}(L,W))\right)^{2}\right]-\mathbb{E}_{P}\left[\left(\partial_{\ell}\widehat{f}(L,W)\right)^{2}\right]$
$=\mathbb{E}_{P}[\rho_{P}(L,W)^{2}(Y-f_{P}(L,W))^{2}]=\mathbb{E}_{P}[\rho_{P}(L,W)^{2}\mathrm{var}_{P}(Y\,|\,L,W)]\geq c^{2}.$
@@
$|f_{P}(1,W_{i})|^{2+\delta}=\frac{|\mathbb{E}_{P}[Y_{i}\mathbbm{1}_{\{L_{i}=1\}}\,|\,W_{i}]|^{2+\delta}}{\pi_{P}(W_{i})^{2+\delta}}\leq\mathbb{E}_{P}\left[|Y_{i}|^{2+\delta}\,|\,W_{i}\right]c^{-(2+\delta)}$
@@
$\kappa_{P}^{2}=E_{P}[S_{i}]^{2}\leq p_{P}\mathbb{E}_{P}\left[(f_{P}(1,W)-f(L,W))^{2}\right]$
@@
    $T:=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-\mathbb{E}_{P}[X])}{\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\right)^{2}\right)^{1/2}}\overset{d/\mathcal{P}}{\rightrightarrows}\mathcal{N}(0,1).$
@@
    $\widetilde{T}:=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-\mathbb{E}_{P}[X])}{\left(\mathrm{var}_{P}(X)\right)^{1/2}}\overset{d/\mathcal{P}}{\rightrightarrows}\mathcal{N}(0,1).$
@@
Let $(L,W)$ denote random variables taking values in ×$\mathbb{R}\times\Delta^{d-2}$ and let $P_{W}$ denote the distribution of $W$. Suppose that the conditional distribution of $L$ given $W=w$ is absolutely continuous with respect to Lebesgue measure with conditional density $p(\ell\,|\,w)$. Let $g:\mathbb{R}\times\Delta^{d-2}\to\mathbb{R}$ denote a function. Assume that for $P_{W}$-almost every $w\in\Delta^{d-2}$:
@@
The following is an adaptation of Klyne and Shah [2023, Proposition 1] (the primary change is the slight rephrasing of assumption (d)). We prove the result by showing that for $P_{W}$-almost all $w\in\Delta^{d-2}$, we have
@@
Multiplying both sides of the inequality by $\mathbbm{1}_{[x,x+1]}(y)$, using monotonicity of integrals and Fubini’s theorem, we see that

### v2

@@ 4.3 Variable influence measures when predicting BMI from gut microbiome
While the average perturbation effect estimates which are specifically tailored to the compositional setting are consistent across different regression methods, the permutation and $R^{2}$ measures are – unsurprisingly – not. The population version of the nonparametric $R^{2}$ is always 00, so the estimates are essentially mean zero random noise that depends on the estimator. For the permutation measure the inconsistency comes from the fact that the permuted data inherently leave the support of the simplex, making the resulting influence measure highly dependent on how the regression methods extrapolate. Unfortunately, due to a lack of alternatives, it is still common to use these types feature importance or influence measures in microbiome science (e.g., Marcos-Zambrano et al., 2021). Furthermore, the results show that the nonparametric estimator of the CKE is less well-behaved than the other estimators based on the partially linear model. One reason for this could be that dividing by the propensity score estimator results in a more unstable estimate, particularly when the propensity score is poorly estimated.
Next, we empirically demonstrate that simply ignoring the simplex constraint is not an option. To this end we consider a standard approach for estimating effects of coordinates for unconstrained Euclidean data. More specifically, we compare the proposed $\theta_{j}=\mathrm{CFI}_{\mathrm{unit}}^{j}$ (with estimator based on the partially linear model) with a similar estimator but based on the partially linear model $\mathbb{E}[Y\,|\,Z]=\theta_{j}Z^{j}+h(Z^{-j})$ in which the parameter $\theta_{j}$ is unidentified due to the simplex constraint. For both methods we use a regression estimator that uses cross-validation to select the best estimator out of a linear regression, random forest and support vector regression. As expected the results from the estimator that ignores the simplex structure are meaningless; for example, $\mathrm{var}(Z^{j})\widehat{\theta}_{j}^{2}/\mathrm{var}(Y)$, a proxy for the variance explained by the linear effect, is larger than $10^{17}$ for all species. In contrast, for $\mathrm{CFI}_{\mathrm{unit}}$ these values lie between $10^{-7}$ and $1.6$.
@@ 6 Discussion
We introduced perturbations on the simplex as a formalization of hypothetical changes to a compositional vector and used them to construct interpretable compositional effect measures, which we call average perturbation effects. In our view, perturbations provide a natural way for practitioners to formalize research questions about non-standard changes to a system. We illustrated this, for example, by using perturbations to formalize the otherwise vague concept of an effect of diversity as well as to construct nonparametric feature influence measures for compositional coordinates. Unfortunately, many applied researchers working with compositional data, generally due to a lack of alternatives, still rely on marginal effects by analyzing dependence between a summary statistic and a response (e.g., by regressing a response on a diversity measure). As we demonstrated, such approaches can be misleading since they do not account for the confounding induced by the remaining variation in the compositional vector. We hope that this work can draw attention to such pitfalls and improve the analysis of such data by providing a set of easily accessible alternative methods.
@@
We additionally proposed efficient estimators for the average perturbation effects based on semiparametric theory and provided theoretical guarantees that enable us to perform valid inference. Most statistical methods for compositional data analysis to date are based on the parametric log-ratio approach introduced in the seminal paper Aitchison (1982) or other parametric and transformation-based approaches. Our nonparametric approach goes beyond these established methods and allows practitioners to apply theoretically justified procedures to modern compositional data applications. Importantly, our target parameters and estimators circumvent the zero problem central to any log-ratio based analysis and naturally apply to high-dimensional data. This makes our proposed estimators applicable to a range of modern applications that previously had to rely on ad hoc methods or simply ignore the compositional constraints. An additional advantage of a perturbation-based analysis is that it easily facilitates a causal rather than observational interpretation, allowing practitioners to make standard causal inference assumptions to derive rigorous causal conclusions.
@@
In this work we have focused on data constraints induced by the simplex. There are however many other types of constraints that can arise, e.g., directional data, general manifold-valued data or otherwise constrained Euclidean data. The idea of defining effect measures via perturbations easily extends to those settings. Similarly, perturbations can also be used to define meaningful effects of summary statistics on unconstrained data; something which has received surprisingly little attention in the literature but relates to problems in causal inference (e.g., how to define causal variables). Nevertheless, for both extensions, more work is required to develop reparametrizations that can be used to derive efficient estimators akin to ours using semiparametric theory. We hope our work will spark new developments in this area and help provide practitioners with the statistical tools they need.
@@ Acknowledgements
We thank Alexander Mangulad Christgau and Lucas Kook for helpful discussions. We also thank Viktor Skjold Lindegaard for spotting some minor errors in an earlier version of the manuscript. ARL and NP were supported by a research grant (0069071) from Novo Nordisk Fonden.
@@
(b) By the same argument as in (a), we have that $\phi$ is invertible with inverse as computed in (S3). This reparametrization is derivative-isolating for perturbations with direction $v_{E}$ given by (6) and speed given by (reversing the computation in (a))
$s:z\mapsto-\partial_{\ell}t_{(E(z),v_{E}(z))}^{-1}(t(z))=-(\partial_{\delta}t_{(E(z),v_{E}(z))}(\|E(z)-z\|_{1}))^{-1}.$
@@
$L:=W^{1}+(1+\mathds{1}_{\{W^{1}>1/2\}})\xi,\quad Y:=(1+L)W^{1}+\varepsilon.$
@@
$\min_{\theta\in\mathbb{R},\widetilde{h}:\mathcal{W}\to\mathbb{R}}\underbrace{\mathbb{E}\left[\left(\mathbb{E}[Y\,|\,L,W]-\mathbb{E}[Y\,|\,W]-\theta(L-\mathbb{E}[L\,|\,W])-\widetilde{h}(W)\right)^{2}\right]}_{=c(\theta,\widetilde{h})},$ (S5)
@@
2:For $i\in[n]$, define $D_{i}:=\mathds{1}_{\{\omega_{\psi}(Z_{i})\neq 0\}}$.
@@
Section S4.1 provides the details behind the experiments in Table 1 in Section 1. Section S4.2 contains additional simulation results from our investigation into the performance of various semiparametric estimators. Section S4.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 4.2.1. Section S4.5 provides a detailed description of the non-compositional background variables that are controlled for in Section 4.2.2.
@@
$Y\,|\,Z\sim\mathrm{Bern}\left(\frac{3}{4}-\frac{1}{8}\mathds{1}_{\{Z^{1}=0\}}+\frac{1}{2}\mathds{1}_{\{Z^{2}=0\}}\right).$
@@
S4.3 Description of the construction of groups in Section 4.2.1
In this section we describe how we construct the two community-level grouped datasets analyzed in Section 4.2.1. On the individual-level data the observed proportions of race (categorical, ’White’, ’Black’ and ’Other’), education (binary, $1$ if strictly greater than high school and 00 otherwise) and compensation (binary, $1$ if compensation above 50,000 USD per year and 00 otherwise) are
@@
S4.4 Semi-synthetic individual-level experiment using groups from Section 4.2.1
In this section we consider a variant of the experiment performed in Section 4.2.1 where instead of aggregating the full dataset into groups, we preserve the individual observations but generate new responses based on the groups. More specifically, we generate the groups as before and estimate the probability that $Y=1$ within each group and use this probability to generate new $Y$s for each observation. For each individual, we thus have a response $Y$, a race $R$, a ‘contextual’ variable $Z$ that gives the racial distribution of each group and additional covariates $X$. By construction of the groups and sampling of the $Y$, we know that there is a positive effect of diversity on the income. We now apply the naive_diversity and $\mathrm{CDI}_{\mathrm{Gini}}$ methods as in Section 4.2.1 except we always control for the race of inviduals and omit the unstable estimators that condition on $Z$. The results can be seen in Table 5.
Method Grouping on compensation
Estimate $95\%$ CI
[…]
$\mathrm{CDI}_{\mathrm{Gini}}\,|\,R$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.459}}$ $(0.247,0.671)$
$\mathrm{CDI}_{\mathrm{Gini}}\,|\,R,X$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.487}}$ $(0.282,0.692)$
Table 5: Table of estimates in the individual-level ‘Adult’ data experiment. Bold red numbers indicate a significant negative effect of diversity at a 5% level, while bold green indicates a significant positive effect. The naive diversity method indicates that there is a significant negative effect of diversity on compensation, even when controlling for individual race ($R$) and additional covariates ($X$), despite the fact that the data is constructed with a positive effect of diversity. The CDI correctly identifies that the effect is positive.
We see that, as in Section 4.2.1, the naive_diversity method incorrectly estimates a negative effect of diversity on income while $\mathrm{CDI}_{\mathrm{Gini}}$ correctly identifies that the effect is positive.
S4.5 Description of additional control variables used in Section 4.2.2
The additional background variables in the New York schools data used in Section 4.2.2 are described below. Unless otherwise mentioned, no preprocessing was used.
School income estimate
The school income estimates were grouped into four categories; one for missing values and three more based on quantiles of the observed values.
School grade range (grade referring to years of school)
The range of grades for the schools were grouped into three categories; one for ‘high grades’ where the lowest grade was greater or equal to 6th, one for ‘low grades’ where the highest grade was less than or equal to 4th and the remaining schools were put in a ‘mixed’ category.
Student attendance rate
This number was computed as the total number of days attended by all students / total number of days on register for all students.
% chronically absent students
The percentage of students missing 10% of school days or at least 18 days per year in a 180-day school year.
Economic Need Index
Computed as the percentage of students in temporary housing plus 0.5 times the percentage of students that are eligible for Health Reimbursement Arrangements plus 0.5 times the percentage of students eligible for free lunch.
Community school
An indicator of whether the school is a community school.
% English language learners
Percentage of students who are learning English for the first time.
Rigorous Instruction %
School rating for ‘how well the curriculum and instruction engage students, build critical-thinking skills, and are aligned to the Common Core’.
Collaborative Teachers %
School rating for ‘how well teachers participate in opportunities to develop, grow, and contribute to the continuous improvement of the school community’.
Supportive Environment %
School rating for ‘how well the school establishes a culture where students feel safe, challenged to grow, and supported to meet high expectations’.
Effective School Leadership %
School rating for ‘how well school leadership inspires the school community with a clear instructional vision and effectively distributes leadership to realize this vision’.
Strong Family-Community Ties %
School rating for ‘how well the school forms effective partnerships with families to improve the school’.
Trust %
School rating for ‘whether the relationships between administrators, educators, students, and families are based on trust and respect’.
S4.6 Additional plots from gut microbiome analysis in Section 4.3
@@
$\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon)\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(\{|(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon\}\cap\Omega_{P})+\zeta/2\leq\sup_{P\in\mathcal{P}}\mathbb{P}_{P}(|\mathds{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon)+\zeta/2,$
hence it suffices to prove that $\mathds{1}_{\Omega_{P}}\cdot(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ to conclude that $(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$. To that end, we first note that
$\mathbb{E}_{P}\left[\mathds{1}_{\Omega_{P}}\cdot(\textup{III})\,|\,\widehat{f}_{k}\right]=\mathbb{E}_{P}\left[\mathds{1}_{\Omega_{P}}\left(\partial_{\ell}\widehat{f}_{k}(L_{i},W_{i})-\partial_{\ell}f_{P}(L_{i},W_{i})+\rho_{P}(L_{i},W_{i})\{\widehat{f}_{k}(L_{i},W_{i})-f_{P}(L_{i},W_{i})\}\right)\,|\,\widehat{f}_{k}\right].$
@@
$\mathbb{E}_{P}\left[\mathds{1}_{\Omega_{P}}\cdot(\textup{III})\,|\,\widehat{f}_{k}=h\right]=0$ (S11)
for $\mathbb{P}_{P}$-almost every $h$. If $h$ is such that $\mathds{1}_{\Omega_{P}}=0$ then clearly the conditional expectation is 00. If $h$ is such that $\mathds{1}_{\Omega_{P}}=1$, then we intend to apply Lemma S5 (integration by parts) with $g=(h-f_{P})$. Assumptions (a)-(c) of Lemma S5 are clear from the above. Assumption (d) of Lemma S5 also holds, since then both $\mathbb{E}[|h(L,W)-f_{P}(L,W)|\,|\,W]$ and $\mathbb{E}[|\partial_{\ell}(h(L,W)-f_{P}(L,W))+\rho_{P}(L,W)(h(L,W)-f_{P}(L,W))|\,|\,W]$ are almost surely finite when $\mathds{1}_{\Omega_{P}}=1$, so Lemma S6 yields that the assumption is satisfied. We conclude that
@@
$\mathbb{P}_{P}(|\mathds{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\epsilon\,|\,\widehat{f}_{k})$
$\qquad\leq\frac{1}{\epsilon^{2}N_{k,r}}\mathbb{E}_{P}[\mathds{1}_{\Omega_{P}}\cdot(\textup{III})^{2}\,|\,\widehat{f}_{k}]$
@@
$\mathbb{P}_{P}(|\mathds{1}_{\Omega_{P}}\cdot(\textup{III})|\geq N_{k,r}^{-1/2}\,|\,\widehat{f}_{k})=o_{\mathcal{P}}(1)$
which implies that $\mathds{1}_{\Omega_{P}}\cdot(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ so $(\textup{III})=o_{\mathcal{P}}(N_{k,r}^{-1/2})$ by our earlier argument. The above combined with Lemma S1 proves (S10).
@@
and (S12) implies that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(S_{i}-\tau_{P})=O_{\mathcal{P}}(1)$, we have by Lemma S3, (S10) and (S13) that
@@
$\mathrm{var}_{P}(S_{i})$ $=\mathbb{E}_{P}\left[\left(\partial_{\ell}f_{P}(L,W)-\rho_{P}(L,W)(Y-f_{P}(L,W))\right)^{2}\right]-\mathbb{E}_{P}\left[\partial_{\ell}f_{P}(L,W)\right]^{2}$
$\geq\mathbb{E}_{P}[\rho_{P}(L,W)^{2}(Y-f_{P}(L,W))^{2}]=\mathbb{E}_{P}[\rho_{P}(L,W)^{2}\mathrm{var}_{P}(Y\,|\,L,W)]\geq c^{2}.$
@@
$|f_{P}(1,W_{i})|^{2+\delta}=\frac{|\mathbb{E}_{P}[Y_{i}\mathds{1}_{\{L_{i}=1\}}\,|\,W_{i}]|^{2+\delta}}{\pi_{P}(W_{i})^{2+\delta}}\leq\mathbb{E}_{P}\left[|Y_{i}|^{2+\delta}\,|\,W_{i}\right]c^{-(2+\delta)}$
@@
$\kappa_{P}^{2}=\mathbb{E}_{P}[S_{i}]^{2}\leq p_{P}\mathbb{E}_{P}\left[(f_{P}(1,W)-f(L,W))^{2}\right]$
@@
    $T:=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-\mathbb{E}_{P}[X])}{\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}^{2}-\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\right)^{2}\right)^{1/2}}\quad\text{satisfies}\quad\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\sup_{t\in\mathbb{R}}|\mathbb{P}_{P}(T\leq t)-\Phi(t)|=0$
@@
    $T:=\widetilde{T}:=\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(X_{i}-\mathbb{E}_{P}[X])}{\left(\mathrm{var}_{P}(X)\right)^{1/2}}\quad\text{satisfies}\quad\lim_{n\to\infty}\sup_{P\in\mathcal{P}}\sup_{t\in\mathbb{R}}|\mathbb{P}_{P}(\widetilde{T}\leq t)-\Phi(t)|=0$
@@
[…]
@@
The following is an adaptation of Klyne and Shah [2023, Proposition 1] (the primary change is the slight rephrasing of assumption (d)). We prove the result by showing that for $P_{W}$-almost all $w\in\mathcal{W}$, we have
@@
Multiplying both sides of the inequality by $\mathds{1}_{[x,x+1]}(y)$, using monotonicity of integrals and Fubini’s theorem, we see that